The perimeter of an equilateral triangle, triangle RST is 72 inches. What is the length of an altitude of triangle RST

A. 24 inches
B. 12 radical 3 inches
C. 12 radical 2 inches
D. 12 inches

Answer by richwmiller(17219)

9*8=72
3*24=72
each side is 24
the base is 12
12^2+b^2=24^2
b=12sqrt(3) which is choice B)